OpenStudy (anonymous):

the weekly sales of a magazine increased from 500.000 to 600,000 by what percentage did the magazine sale increase? this kind of hard

rebeccaxhawaii (rebeccaxhawaii):

what is 600,000-500,000=?

OpenStudy (anonymous):

i dont no u tell me

OpenStudy (anonymous):

600.000-500,000=1000 i guess

OpenStudy (anonymous):

100,000/500,000 x 100 =100/5 =20%

rebeccaxhawaii (rebeccaxhawaii):

nope you keep missing the zeros

OpenStudy (anonymous):

how did u get 20% can u draw for me please

OpenStudy (anonymous):

100,000 Came because 600,000- 500,000=100,000

rebeccaxhawaii (rebeccaxhawaii):

@thevi please dont give out answers

OpenStudy (anonymous):

Formula is Increase of prices which is 100,000 over 500,000 into 100 The way in which u find the percentage for any fraction is by multiplying by 100

OpenStudy (anonymous):

did u multiplaying 100.000X 500.000

OpenStudy (anonymous):

Sorry I am using the mobile app so I can't draw 100 is 100/1 it is multiplication of fractions

OpenStudy (anonymous):

(100,000/500,000) x (100/1)

OpenStudy (anonymous):

this kind of hard for me :(

OpenStudy (anonymous):

i didnt get answer

OpenStudy (anonymous):

Did it help u

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

In 100,000/500,000 fraction you can simplify it by cutting off the zeros Then u get 1/5 So (1/5) x (100/1) = 100/5 U can simplify this by dividing both sides by 5 So u end up with 20/1 Therefore your answer is 20%